The difference of two numbers is 2, and their sum is 4. What are the two numbers?

Answer by TimothyLamb(4379)

n - m = 2
n = 2 + m
m + n = 4
---
m + n = 4
m + 2 + m = 4
2m = 2
---
m = 1
n = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php